determine the slope and the y-intercept. then graph the equation. y=2/3x+5
Answer by stanbon(75887)

determine the slope and the y-intercept. then graph the equation. y=2/3x+5
-----
slope = 2/3
intercept = 5
-----
Put your pencil point at (0,5) on the y-axis.
----
Move your pencil point 3 to the right to (3,5)
Then move your pencil point (2/3)3 = 2 units up to (3,7)
---
You have points at (0,5) and (3,7). Draw a line thru them.
----
That is the graph you want.
============
Cheers,
Stan H.
=================